A rectangle has an area of 36cm squared. One dimension is odd. The odd dimension is also a prime number. One dimension is even. Then find the perimeter of the rectangle. Respond fully then explain your work.
Answer by mananth(16946)

Area = 36
Area = L * W
Find factors of 36
2*2*3*3
The options for the lengths are
2*2=4 &3*3=9 (4,9)
2*2*3=12 &3 (12,3)
2*3=6 & 2*3=6 (6,6)
3*3*2=18 &2 (18,2)
Pick out the sides for the condtion
it is 12 cms*3cm


One is odd & prime and other is even